# POINTS: Improving Your Vision-language Model with Affordable Strategies

### v2

@@ 1 Introduction
Advancements in large language models (LLMs; Chowdhery et al. 2023, Jiang et al. 2023, OpenAI 2022, Yang et al. 2024, Dubey et al. 2024) have significantly enhanced the capabilities of vision-language large models (Fu et al. 2023, Liu et al. 2023b, OpenAI 2023, Dong et al. 2024a, Zhu et al. 2023), enabling more sophisticated analyses of textual and visual information. Prominent closed-source model paradigms such as GPT-4 (OpenAI, 2023), Gemini Pro 1.5 (Fu et al., 2023), and Claude 3 (Anthropic, 2024) have achieved remarkable success in expanding LLMs into the realm of vision-language models. Concurrently, open-source vision-language large models are also advancing rapidly, with numerous notable contributions emerging in the field (Liu et al., 2024b; Chen et al., 2024c).
@@
Historically, LLaVA (Liu et al., 2024b) has served as a common baseline. However, recent advancements have rendered its performance suboptimal. Thus, there is a need to establish a stronger baseline for further exploration. In this work, we enhance the vanilla LLaVA architecture by refining the pre-training dataset. Inspired by CapFusion (Yu et al., 2024), we merge the original captions with world knowledge and generated captions that exhibit good grammatical structure. For visual instruction tuning datasets, we introduce Individual Select (Liu et al., 2024c) to curate effective instruction tuning datasets. Regarding model architecture, we first incorporate Dynamic High Resolution to help the model capture fine-grained details. To address image distortion issues inherent in Dynamic High Resolution, we propose a novel image splitting strategy called Consistent Aspect Ratio Dynamic High Resolution, which maintains a consistent image ratio. Additionally, inspired by Vary (Wei et al., 2023), we merge features from a vision encoder trained separately with text-rich data with those from the original vision encoder, significantly boosting the model’s OCR capabilities. Unlike most existing works (Li et al., 2024; Chen et al., 2024c), we extensively ablate each newly introduced component in the strong baseline to verify their individual benefits.
@@
Recent works seldom explore the optimization of pre-training datasets. Most studies (Chen et al., 2024c; Yao et al., 2024; Bai et al., 2023b) tend to empirically combine samples from various large-scale datasets (Schuhmann et al., 2022; Byeon et al., 2022), often leading to inefficient and computationally expensive pre-training processes. In the domain of large language models, some research leverages perplexity to filter pre-training datasets. Inspired by this approach, we filter our pre-training dataset by selecting the top samples with the lowest perplexity values. This filtering process yields a subset of 1 million data samples, on which we subsequently pre-train our model. Experimental results demonstrate that the model trained on this filtered subset outperforms a model trained on a dataset five times larger.
@@
In the visual instruction tuning stage, most existing works (Liu et al., 2024c; Li et al., 2024; Chen et al., 2024c) focus on collecting large quantities of datasets and performing ablation studies to select the most effective ones. However, this approach often reaches a plateau, where introducing additional datasets yields only marginal or even degraded performance. Previous research on model soup has demonstrated the benefits of merging weights from different models fine-tuned with various hyper-parameters. In this work, we propose using model soup to merge weights from models fine-tuned with different datasets to further improve performance when dataset selection no longer brings significant improvement. Compared to conducting model soup on models fine-tuned with different hyper-parameters, e.g. learning rate, the improvement with model soup on models fine-tuned with different datasets is much more prominent. Following this line of work, we further experiment with different model soup strategies and find that greedy model soup is the most effective.
@@ 2 Related Works
The rapid advancement of large language models (LLMs; Dubey et al. 2024, Team et al. 2023, Achiam et al. 2023, Yang et al. 2024, Su et al. 2022) has laid the groundwork for the emergence of multimodal large language models (MLLMs; Li et al. 2024, Liu et al. 2024b, Liu et al. 2024a, Bai et al. 2023a, Qiao et al. 2024), which aim to integrate visual understanding with language reasoning and multimodal perception and comprehension. Prominent models such as GPT-4v (Achiam et al., 2023) and Gemini-1.5-Pro (Team et al., 2023), developed by major corporations, have spearheaded the MLLM era, utilizing proprietary training data and undisclosed training methodologies. Meanwhile, open-source models have been striving to keep pace. For instance, LLaVA-Next (Liu et al., 2024a) and InternVL-1.5 (Chen et al., 2024c) introduce dynamic high-resolution techniques by dividing a large image into multiple smaller segments with ratio-inconsistent resizing. MiniCPM-V (Yao et al., 2024) employs a specialized vision encoder to generate non-square image patches. Additionally, models like Vary (Wei et al., 2023), SPHINX (Lin et al., 2023), Cambrian-1 (Tong et al., 2024), and Mini-Gemini (Li et al., 2023) propose dual vision encoders to enhance visual capabilities. Furthermore, the significant progress in multimodal model evaluation (Liu et al., 2023c; Chen et al., 2024b; Fang et al., 2024) has also contributed to the rapid improvement of large vision-language models. In this work, we introduce POINT, a model trained exclusively with fully open-source datasets during both the pre-training and supervised fine-tuning (SFT) stages, demonstrating promising results on extensive benchmarks.
@@
The selection of training data for multimodal models is of paramount importance (Laurençon et al., 2024; Tong et al., 2024), and most improvements in existing works stem from detailed ablation of instruction tuning datasets (Li et al., 2024; Liu et al., 2024b; Chen et al., 2024c). The commonly used approach to select the most effective datasets involves iteratively adding each dataset to the pool; if it brings improvement, we keep it, otherwise, we drop it. However, this approach may eventually plateau, as further additions might only yield marginal improvements. To further enhance performance, we propose employing model soup (Wortsman et al., 2022) on different models fine-tuned with various visual instruction tuning datasets. This method involves merging the model weights after visual instruction tuning on diverse datasets, resulting in notable performance improvements.
@@ 3 Methods
This section is divided into three parts: i) In subsection 3.1, we integrate various techniques from previous methods (Liu et al., 2024a; Lin et al., 2023; Wei et al., 2023; Liu et al., 2024c; Chen et al., 2024c) to create a strong baseline for further experiments. Additionally, we propose a novel dynamic resolution splitting method, termed Consistent Aspect Ratio Dynamic High Resolution (CATTY for short), to mitigate the issue of image distortion. ii) In subsection 3.2, we propose using perplexity to filter the pre-training dataset. iii) Finally, in subsection 3.3, we incorporate the concept of model soup (Wortsman et al., 2022) into the instruction tuning stage. We find that this straightforward approach can significantly improve the model’s performance, especially when further data selection only brings marginal or even degraded performance.
@@ 3.1 A Strong Baseline
In this section, we integrate the recent advancements from existing works to create a strong baseline, containing Dynamic High Resolution from LLaVA-Next(Liu et al., 2024a) and InternVL1.5(Chen et al., 2024c), CapFusion from (Yu et al., 2024), Dual Vision Encoder from Vary(Wei et al., 2023) and SPHINX(Lin et al., 2023), Individual Select from (Liu et al., 2024c). Following LLaVA(Liu et al., 2024b), POINTS mainly contains three parts: vision encoder, projector and the large language model. By integrating all these practices from previous works, we obtain the model structure and pipeline in Figure 1.
@@
It has been verified that feeding high-resolution images to vision-language models is beneficial for capturing fine-grained details and reducing hallucinations (Liu et al., 2023b). To enable vision encoder with fixed input resolutions to accommodate dynamic image resolutions, Dynamic High Resolution in LLaVA-Next (Liu et al., 2024a) and InternVL-1.5 (Chen et al., 2024c) splits high-resolution images into several tiles of the same resolution, which the original vision encoder can process. The concrete steps are as follows: i) First, the maximum number of tiles an image can be split into is predefined (set to 8 in our experiments). ii) Based on the maximum number of tiles, a table is created containing information about the target image before splitting. The key of the table is the aspect ratio, and the value is the width and height of the target image, which can be evenly divided by the resolution of the vision encoder. iii) For each image, the target resolution is fetched from the pre-computed table according to the similarity between aspect ratios. The current image is then resized to the target resolution and split into several tiles of the same resolution.
@@
Before splitting the image, Dynamic High Resolution in InternVL-1.5 (Chen et al., 2024c) resizes the image to the target resolution. However, this resizing is not proportional to the image’s original aspect ratio, which can cause distortion. This issue has been discussed in previous articles(Yao et al., 2024). Therefore, we propose a splitting method that maintains the image’s aspect ratio, named Consistent Aspect Ratio Dynamic High Resolution (see Figure 2). The first two steps in CATTY are the same as those in InternVL-1.5, and the last step works as follows: Given an image with height $\mathrm{H}$ and width $\mathrm{W}$, we obtain the height and width of the referenced image from the pre-computed table, denoted as $\mathrm{H}^{\text{r}}$ and $\mathrm{W}^{\text{r}}$, respectively. Then, we resize the image to the target size ($\mathrm{H}^{\text{t}}\times\mathrm{W}^{\text{t}}$) by:
@@
In Equation 2, $\mathrm{S}^{\text{h}}$ is set to 0 if $\mathrm{H}^{\text{r}}/\mathrm{H}^{\text{v}}=1$, and similarly for $\mathrm{S}^{\text{w}}$. This approach allows us to divide a high-resolution image into several tiles without introducing any distortion. Alongside the tiles obtained using CATTY, we also include a thumbnail of the global view of the image to capture the overall context. This thumbnail is resized to match the input resolution of the vision encoder. Before feeding the features output by the vision encoder into the large language model, we employ the pixel shuffle technique with a down-sampling factor of 0.25, as described in InternLM-XComposer2-4KHD (Dong et al., 2024b), to reduce the sequence length of the image features for improved efficiency.
@@
Several previous works, such as SPHINX (Lin et al., 2023) and Cambrian-1 (Tong et al., 2024), have demonstrated that different vision encoders exhibit distinct advantages across various domains. Combining features from multiple encoders can lead to improved and more robust performance. Unlike the perception and reasoning required for natural images, text-intensive images demand different capabilities from vision-language models (Wei et al., 2023). To enhance optical character recognition (OCR) capabilities, we train a separate vision encoder, referred to as the OCR ViT, to extract textual features from images, following the methodology of Vary (Wei et al., 2023). Unlike Vary, we do not construct training samples, such as charts, ourselves; instead, we utilize OCR results (extracted using PaddleOCR in our case) for pre-training. Additionally, we include natural captions in the pre-training dataset for the OCR vision encoder. More details about the composition of the pre-training datasets for the OCR vision encoder will be discussed in the following section. We merge the features from the general vision encoder (called General ViT) and the OCR vision encoder using a weighted average before feeding them into the large language model.
@@
Individual Select, as proposed by Liu et al. (2024c), aims to identify the most effective instruction tuning datasets. Building on this approach, we adopt the dataset composition from Liu et al. (2024c) as our candidate pool and incorporate additional datasets used in DeepSeek-VL (Lu et al., 2024a), Cambrian-1 (Tong et al., 2024), and Cauldron (Laurençon et al., 2024). Ultimately, we integrate 16 more datasets into those identified by Liu et al. (2024c) (further details are provided in subsection 4.2). To enhance the diversity of prompts, given the homogeneity in the style of prompts within academic datasets, we employ GPT-4o to generate question-answer pairs in line with previous works (Lu et al., 2024a; Chen et al., 2024c) (the prompt to generate question-answer pairs will be provided in the appendix). The images for these pairs are randomly selected from LAION-5B (Schuhmann et al., 2022). We refer to the final composition of visual instruction tuning datasets as the Base Set.
@@ 3.3 Instruction Data Selection with Model Soup
Visual instruction tuning data is crucial for the superior performance of existing vision-language models (Chen et al., 2024c; Dong et al., 2024a; Liu et al., 2024b). However, most existing works focus on selecting more effective datasets by iterative ablation. In many cases, this approach reaches a plateau, where further data selection can only bring marginal improvements or even degrade performance. In this section, we introduce the benefits of using model soup to integrate the advantages of models fine-tuned with different instruction tuning datasets after data selection meets a bottleneck. The philosophy behind model soup is as follows: given a pre-trained model, fine-tuning the model with different hyper-parameters, $h_{1},\ldots,h_{k}$, results in several fine-tuned models converging to different local optima, denoted as $f(\theta_{1},h_{1}),\ldots,f(\theta_{k},h_{k})$. These hyper-parameters include learning rate, data augmentation, initialization seed, etc. By interpolating the weights of these fine-tuned models, we can always obtain a stronger model, $f(\theta_{s},h_{s})$. Given the pre-trained model obtained through the methods discussed above, a base instruction tuning dataset $D$, and a series of visual instruction tuning datasets $d_{1},\ldots,d_{k}$ to be selected, we can obtain a stronger model using the following steps:

### v3

@@ 1 Introduction
Advancements in large language models (LLMs; Chowdhery et al. 2023, Jiang et al. 2023, OpenAI 2022, Yang et al. 2024, Dubey et al. 2024) have significantly enhanced the capabilities of vision-language large models (Fu et al. 2023, Liu et al. 2023b, OpenAI 2023, Dong et al. 2024a, Zhu et al. 2023), enabling more sophisticated analyses of textual and visual information. Prominent closed-source model paradigms such as GPT-4 (OpenAI, 2023), Gemini Pro 1.5 (Fu et al., 2023), and Claude 3 (Anthropic, 2024) have achieved remarkable success in expanding LLMs into the realm of vision-language models. Concurrently, open-source vision-language large models are also advancing rapidly, with numerous notable contributions emerging in the field (Liu et al., 2024b; Chen et al., 2024d).
@@
Historically, LLaVA (Liu et al., 2024b) has served as a common baseline. However, recent advancements have rendered its performance suboptimal. Thus, there is a need to establish a stronger baseline for further exploration. In this work, we enhance the vanilla LLaVA architecture by refining the pre-training dataset. Inspired by CapFusion (Yu et al., 2024), we merge the original captions with world knowledge and generated captions that exhibit good grammatical structure. For visual instruction tuning datasets, we introduce Individual Select (Liu et al., 2024c) to curate effective instruction tuning datasets. Regarding model architecture, we first incorporate Dynamic High Resolution to help the model capture fine-grained details. To address image distortion issues inherent in Dynamic High Resolution, we propose a novel image splitting strategy called Consistent Aspect Ratio Dynamic High Resolution (CATTY), which maintains a consistent image ratio. Additionally, inspired by Vary (Wei et al., 2023), we merge features from a vision encoder trained separately with text-rich data with those from the original vision encoder, significantly boosting the model’s OCR capabilities. Unlike most existing works (Li et al., 2024a; Chen et al., 2024d), we extensively ablate each newly introduced component in the strong baseline to verify their individual benefits.
@@
Recent works seldom explore the optimization of pre-training datasets. Most studies (Chen et al., 2024d; Yao et al., 2024; Bai et al., 2023b) tend to empirically combine samples from various large-scale datasets (Schuhmann et al., 2022; Byeon et al., 2022), often leading to inefficient and computationally expensive pre-training processes. In the domain of large language models, some research leverages perplexity to filter pre-training datasets. Inspired by this approach, we filter our pre-training dataset by selecting the top samples with the lowest perplexity values. This filtering process yields a subset of 1 million data samples, on which we subsequently pre-train our model. Experimental results demonstrate that the model trained on this filtered subset outperforms a model trained on a dataset five times larger.
@@
In the visual instruction tuning stage, most existing works (Liu et al., 2024c; Li et al., 2024a; Chen et al., 2024d) focus on collecting large quantities of datasets and performing ablation studies to select the most effective ones. However, this approach often reaches a plateau, where introducing additional datasets yields only marginal or even degraded performance. Previous research on model soup has demonstrated the benefits of merging weights from different models fine-tuned with various hyper-parameters. In this work, we propose using model soup to merge weights from models fine-tuned with different datasets to further improve performance when dataset selection no longer brings significant improvement. Compared to conducting model soup on models fine-tuned with different hyper-parameters, e.g. learning rate, the improvement with model soup on models fine-tuned with different datasets is much more prominent. Following this line of work, we further experiment with different model soup strategies and find that greedy model soup is the most effective.
@@ 2 Related Works
The rapid advancement of large language models (LLMs; Dubey et al. 2024, Team et al. 2023, Achiam et al. 2023, Yang et al. 2024, Su et al. 2022) has laid the groundwork for the emergence of multimodal large language models (MLLMs; Li et al. 2024a, Liu et al. 2024b, Liu et al. 2024a, Bai et al. 2023a, Qiao et al. 2024), which aim to integrate visual understanding with language reasoning and multimodal perception and comprehension. Prominent models such as GPT-4v (Achiam et al., 2023) and Gemini-1.5-Pro (Team et al., 2023), developed by major corporations, have spearheaded the MLLM era, utilizing proprietary training data and undisclosed training methodologies. Meanwhile, open-source models have been striving to keep pace. For instance, LLaVA-Next (Liu et al., 2024a) and InternVL-1.5 (Chen et al., 2024d) introduce dynamic high-resolution techniques by dividing a large image into multiple smaller segments with ratio-inconsistent resizing. MiniCPM-V (Yao et al., 2024) employs a specialized vision encoder to generate non-square image patches. Additionally, models like Vary (Wei et al., 2023), SPHINX (Lin et al., 2023), Cambrian-1 (Tong et al., 2024), and Mini-Gemini (Li et al., 2023a) propose dual vision encoders to enhance visual capabilities. Furthermore, the significant progress in multimodal model evaluation (Liu et al., 2023c; Chen et al., 2024c; Fang et al., 2024) has also contributed to the rapid improvement of large vision-language models. In this work, we introduce POINT, a model trained exclusively with fully open-source datasets during both the pre-training and supervised fine-tuning (SFT) stages, demonstrating promising results on extensive benchmarks.
@@
The selection of training data for multimodal models is of paramount importance (Laurençon et al., 2024; Tong et al., 2024), and most improvements in existing works stem from detailed ablation of instruction tuning datasets (Li et al., 2024a; Liu et al., 2024b; Chen et al., 2024d). The commonly used approach to select the most effective datasets involves iteratively adding each dataset to the pool; if it brings improvement, we keep it, otherwise, we drop it. However, this approach may eventually plateau, as further additions might only yield marginal improvements. Previous workHe et al. (2024) has shown the benefits of weight merging, but their experimental results are relatively preliminary. To further enhance performance, we systematically propose employing model soup (Wortsman et al., 2022) on different models fine-tuned with various visual instruction tuning datasets. This method involves merging the model weights after visual instruction tuning on diverse datasets, resulting in notable performance improvements.
@@ 3 Methods
This section is divided into three parts: i) In subsection 3.1, we integrate various techniques from previous methods (Liu et al., 2024a; Lin et al., 2023; Wei et al., 2023; Liu et al., 2024c; Chen et al., 2024d) to create a strong baseline for further experiments. Additionally, we propose a novel dynamic resolution splitting method, termed Consistent Aspect Ratio Dynamic High Resolution (CATTY for short), to mitigate the issue of image distortion. ii) In subsection 3.2, we propose using perplexity to filter the pre-training dataset. iii) Finally, in subsection 3.3, we incorporate the concept of model soup (Wortsman et al., 2022) into the instruction tuning stage. We find that this straightforward approach can significantly improve the model’s performance, especially when further data selection only brings marginal or even degraded performance.
@@ 3.1 A Strong Baseline
In this section, we integrate the recent advancements from existing works to create a strong baseline, containing Dynamic High Resolution from InternVL1.5(Chen et al., 2024d), CapFusion from (Yu et al., 2024), Dual Vision Encoder from Vary(Wei et al., 2023) and SPHINX(Lin et al., 2023), Individual Select from (Liu et al., 2024c). Following LLaVA(Liu et al., 2024b), POINTS mainly contains three parts: vision encoder, projector and the large language model. By integrating all these practices from previous works, we obtain the model structure and pipeline in Figure 1.
@@
It has been verified that feeding high-resolution images to vision-language models is beneficial for capturing fine-grained details and reducing hallucinations (Liu et al., 2023b). To enable vision encoder with fixed input resolutions to accommodate dynamic image resolutions, Dynamic High Resolution in LLaVA-Next (Liu et al., 2024a) and InternVL-1.5 (Chen et al., 2024d) splits high-resolution images into several tiles of the same resolution, which the original vision encoder can process. The concrete steps are as follows: i) First, the maximum number of tiles an image can be split into is predefined (set to 8 in our experiments). ii) Based on the maximum number of tiles, a table is created containing information about the target image before splitting. The key of the table is the aspect ratio, and the value is the width and height of the target image, which can be evenly divided by the resolution of the vision encoder. iii) For each image, the target resolution is fetched from the pre-computed table according to the similarity between aspect ratios. The current image is then resized to the target resolution and split into several tiles of the same resolution.
@@
Before splitting the image, Dynamic High Resolution in InternVL-1.5 (Chen et al., 2024d) resizes the image to the target resolution. However, this resizing is not proportional to the image’s original aspect ratio, which can cause distortion. This issue has been discussed in previous articles(Yao et al., 2024). Therefore, we propose a splitting method that maintains the image’s aspect ratio, named Consistent Aspect Ratio Dynamic High Resolution (see Figure 2). The first two steps in CATTY are the same as those in InternVL-1.5, and the last step works as follows: Given an image with height $\mathrm{H}$ and width $\mathrm{W}$, we obtain the height and width of the referenced image from the pre-computed table, denoted as $\mathrm{H}^{\text{r}}$ and $\mathrm{W}^{\text{r}}$, respectively. Then, we resize the image to the target size ($\mathrm{H}^{\text{t}}\times\mathrm{W}^{\text{t}}$) by:
@@
In Equation 2, $\mathrm{S}^{\text{h}}$ is set to 0 if $\mathrm{H}^{\text{r}}/\mathrm{H}^{\text{v}}=1$, and similarly for $\mathrm{S}^{\text{w}}$. This approach allows us to divide a high-resolution image into several tiles without introducing any distortion. There is one exception: if the aspect ratio of the original image is larger than 8, we resize it to an aspect ratio of 1:8 by default. Alongside the tiles obtained using CATTY, we also include a thumbnail of the global view of the image to capture the overall context. This thumbnail is resized to match the input resolution of the vision encoder. Before feeding the features output by the vision encoder into the large language model, we employ the pixel shuffle technique with a down-sampling factor of 0.25, as described in InternLM-XComposer2-4KHD (Dong et al., 2024b), to reduce the sequence length of the image features for improved efficiency.
@@
Several previous works, such as SPHINX (Lin et al., 2023) and Cambrian-1 (Tong et al., 2024), have demonstrated that different vision encoders exhibit distinct advantages across various domains. Combining features from multiple encoders can lead to improved and more robust performance. Unlike the perception and reasoning required for natural images, text-intensive images demand different capabilities from vision-language models (Wei et al., 2023). To enhance optical character recognition (OCR) capabilities, we train a separate vision encoder, referred to as the OCR ViT, to extract textual features from images, following the methodology of Vary (Wei et al., 2023). Unlike Vary, we do not construct training samples, such as charts, ourselves; instead, we utilize OCR results (extracted using PaddleOCR in our case) for pre-training. Additionally, we include natural captions in the pre-training dataset for the OCR vision encoder. More details about the composition of the pre-training datasets for the OCR vision encoder will be discussed in the following section. We merge the features from the general vision encoder (General ViT) and the OCR vision encoder using a weighted average before feeding them into the large language model.
@@
Individual Select, as proposed by Liu et al. (2024c), aims to identify the most effective instruction tuning datasets. Building on this approach, we adopt the dataset composition from Liu et al. (2024c) as our candidate pool and incorporate additional datasets used in DeepSeek-VL (Lu et al., 2024a), Cambrian-1 (Tong et al., 2024), and Cauldron (Laurençon et al., 2024b). Ultimately, we integrate 16 more datasets into those identified by Liu et al. (2024c) (further details are provided in Appendix). To enhance the diversity of prompts, given the homogeneity in the style of prompts within academic datasets, we employ GPT-4o to generate question-answer pairs in line with previous works (Lu et al., 2024a; Chen et al., 2024d) (the prompt to generate question-answer pairs will be provided in the Appendix). The images for these pairs are randomly selected from LAION-5B (Schuhmann et al., 2022). We refer to the final composition of visual instruction tuning datasets as the Base Set.
@@ 3.3 Instruction Data Selection with Model Soup
Visual instruction tuning data is crucial for the superior performance of existing vision-language models (Chen et al., 2024d; Dong et al., 2024a; Liu et al., 2024b). However, most existing works focus on selecting more effective datasets by iterative ablation. In many cases, this approach reaches a plateau, where further data selection can only bring marginal improvements or even degrade performance. In this section, we systematically introduce the benefits of using model soup to integrate the advantages of models fine-tuned with different instruction tuning datasets after data selection meets a bottleneck. The philosophy behind model soup is as follows: given a pre-trained model, fine-tuning the model with different hyper-parameters, $h_{1},\ldots,h_{k}$, results in several fine-tuned models converging to different local optima, denoted as $f(\theta_{1},h_{1}),\ldots,f(\theta_{k},h_{k})$. These hyper-parameters include learning rate, data augmentation, initialization seed, etc. By interpolating the weights of these fine-tuned models, we can always obtain a stronger model, $f(\theta_{s},h_{s})$. Given the pre-trained model obtained through the methods discussed above, a base instruction tuning dataset $D$, and a series of visual instruction tuning datasets $d_{1},\ldots,d_{k}$ to be selected, we can obtain a stronger model using the following steps: